# Comparison of Extended Lubrication Theories for Stokes Flow

### v1

@@ Appendix D The Backward Facing Step
Due to the BFS having piecewise zero surface gradient, the solutions according to the Reynolds equation, ELT and $\varepsilon^{k}$-PLT models are all equivalent. The pressure contours of this solution are purely one dimensional, and the velocity streamlines are everywhere parallel to $y=0$ instead of wrapping around the sharp corners of the step. In contrast, the Stokes solution is able to capture the discontinuity in the height, leading to a similar flow profile as in the logistic step at large $\lambda$.
The Stokes solutions of pressure and velocity for the BFS and the logistic step at $\lambda=32$ are shown in Fig. 8. The two configurations have the same $H_{\text{in}}/H_{\text{out}}=2$, $L=16$, $\nu=1$, $\mathcal{Q}=1$ and $\mathcal{U}=0$. According to the Stokes solution, the BFS and the logistic step at large $\lambda$ have similar total pressure drops and flow structures. In particular, flow recirculation is significant in the logistic step for $\lambda>16$, with similar positioning to the recirculation observed in the BFS. Yet, as seen in Fig. 2, the Reynolds solution does not exhibit flow recirculation in the logistic step for any $\lambda$, and the VA-ELT and $\varepsilon^{k}$-PLT solutions exhibit flow recirculation at too small of $\lambda$ and with too fast of velocities compared with the Stokes solution.

### v2

@@ Appendix D The Backward Facing Step
Due to the jump discontinuity in $h(x)$ at $x=L/2$, the ELT and $\varepsilon^{k}$-PLT solutions are not well defined. Since the height gradient is piecewise zero, the solution to the Reynolds equation has pressure contours that are purely one dimensional, and the velocity streamlines are everywhere parallel to $y=0$ instead of navigating around the sharp corners of the step. In contrast, the Stokes solution is able to capture the discontinuity in the height, leading to a similar flow profile as in the logistic step at large $\lambda$.
The Stokes solutions of pressure and velocity for the BFS and the logistic step at $\lambda=32$ are shown in Fig. 8. The two configurations have the same $H_{\text{in}}/H_{\text{out}}=2$, $L=16$, $\nu=1$, $\mathcal{Q}=1$ and $\mathcal{U}=0$. According to the Stokes solution, the BFS and the logistic step at large $\lambda$ have similar average pressure drops and flow structures. In particular, flow recirculation is significant in the logistic step for $\lambda>16$, with similar positioning to the recirculation observed in the BFS. Yet, as seen in Fig. 2, the Reynolds solution does not exhibit flow recirculation in the logistic step for any $\lambda$, and the VA-ELT and $\varepsilon^{k}$-PLT solutions exhibit flow recirculation at too small of $\lambda$ and with too fast of velocities compared with the Stokes solution.